OpenStudy (buttermequeasy):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = 1/3x - 3. y = 3x - 14 y = 1/3x + 2 y = -3x + 22 y = -1/3x + 8

OpenStudy (buttermequeasy):

Help please!

OpenStudy (anonymous):

what is the slope of y = 1/3x - 3?

OpenStudy (buttermequeasy):

the slope is 1/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and parallel lines have the same slope so out new line will have that slope too

OpenStudy (buttermequeasy):

Okay. So then it's between the 2 that have 1/3 as the slope

OpenStudy (anonymous):

no, only one has the 1/3 slope

OpenStudy (buttermequeasy):

Oops sorry. I missed the negative sign on the other one.

OpenStudy (anonymous):

want to solve it the rest of the way? or is the answer good enough?

OpenStudy (anonymous):

do you care where the + 2 came from?

OpenStudy (buttermequeasy):

what do you mean do I care?

OpenStudy (anonymous):

do you want to figure it out? does it matter? eventually you will be asked to do it. some problem set, some exam, some time

OpenStudy (buttermequeasy):

Yes.

OpenStudy (anonymous):

ok... so far all we know is the slope of the new line, so y = Mx + b and M=1/3 to find b, you plug in the x and y values from the point given can you do that?